# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ 4.1 Spectral loss functions and approximation classes
    An approximation $\mathcal{C}^{\textup{opt}}_{r}\in\mathscr{C}_{r}$ solves 4.3 if and only if $(\mathcal{C}^{\textup{opt}}_{r})^{-1}\in\mathscr{P}_{r}$ solves 4.4. Furthermore, $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{r}^{\textup{opt}})=\mathcal{L}(\mathcal{C}_{\textup{pos}}\|(\mathcal{P}_{r}^{\textup{opt}})^{-1})$.
Finally, we show that the approximation of $\mathcal{C}_{\textup{pos}}$ or $\mathcal{C}_{\textup{pos}}^{-1}$ by elements of $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$ respectively results in approximate posterior measures that are equivalent to the exact posterior, when keeping the mean fixed. We also shows that 4.3 is well-defined in the sense that $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\cdot)<\infty$ is finite on $\mathscr{C}_{r}$ for any $\mathcal{L}\in\mathscr{L}$. By Lemma 4.5(ii), it follows that 4.4 is analogously well-defined. Recall the definition of $\mathcal{E}$ in (7).
@@
It holds that $\mathcal{C}_{r}\subset\mathcal{E}$. Thus, for any $\mathcal{L}\in\mathscr{L}$, the map $\mathcal{C}\mapsto\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{C})$ is finite on $\mathscr{C}_{r}$ and the map $\mathcal{P}\mapsto\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{P}^{-1})$ is finite on $\mathscr{P}_{r}$.
4.2 Differentiability and minimisers of covariance loss function
In this section we solve 4.4, and by virtue of Lemma 4.5(ii) then also 4.3, by formulating it as a minimisation problem over the set of $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. We want to find the minimiser of the function
$J_{f}:\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\rightarrow\mathbb{R},\quad U\mapsto L_{f}(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}),$
for any $f\in\mathscr{F}$, which we shall express as a composition of functions. This composition will facilitate the analysis of its differentiability and thereby the identification of its stationary points.
Let $\Lambda:L_{2}(\mathcal{H})_{\mathbb{R}}\rightarrow\ell^{2}(\mathbb{R})$ be a function that sends a self-adjoint Hilbert–Schmidt operator to its square summable eigenvalue sequence. We recall that $\ell^{2}(\mathbb{R})$ denotes the space of square integrable sequences. One ordering could be the one that labels the negative eigenvalues with the even integers and the positive eigenvalues with the odd integers, both ordered decreasingly in absolute value. A different choice is to order the eigenvalues in order of decreasing absolute value. We note that the choice of ordering of $\Lambda$ in two operators $T,S\in L_{2}(\mathcal{H})_{\mathbb{R}}$ is allowed to be different. The precise ordering of eigenvalues that $\Lambda$ assigns to an operator is not important, as we shall only look at compositions of $\Lambda$ with a function on $\ell^{2}(\mathbb{R})$ that is permutation invariant.
The function that we shall apply to the image of $\Lambda$ is
which is well-defined by Lemma 4.1(i). Note that if $P\in\mathcal{B}(\ell_{2}((-1,\infty))$ is a permutation, i.e. $((Px)_{i})_{i}=(x_{\alpha(i)})_{i}$ for some bijection $\alpha$ of $\mathbb{N}$, then for every $x\in\ell^{2}((-1,\infty))$ it holds that $F_{f}(Px)=F_{f}(x)$.
$g(U)=\left((\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2}\right)^{*}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I.$ (17)
Then the image of $g$ in fact consists of Hilbert–Schmidt operators and they can be diagonalised in the Cameron–Martin space, as is shown next. This result also motives the definition of $g$, as it shows that $g(U)$ has the same eigenvalues as $R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})$. By eq. 10b, it also shows that $g(U)$ is a nonnegative, self–adjoint update of rank-$r$ of the negative of the posterior-preconditioned Hessian.
Lemma 4.7.
Let $r\in\mathbb{N}$, $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ and $g$ as in (17). Then $\operatorname{rank}\left({g(U)}\right)\leq r+\operatorname{rank}\left({H}\right)$ and there exists a sequence $(e_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ which forms an ONB of $\mathcal{H}$ and a sequence $(\gamma_{i})_{i}\in\ell^{2}((-1,\infty))$ satisfying $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$. Furthermore, it holds that
$g(U)=(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I+\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}.$ (18)
Finally, the eigenvalues of $g(U)$ and $R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})$ agree, counting multiplicities.
Remark 4.8.
In the finite-dimensional setting, $\operatorname{ran}{C}_{\textup{pr}}^{1/2}=\mathcal{H}$ and thus one can simply define $g(U)$ on all of $\mathcal{H}$ as
$\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}-I$ $=\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1/2}(I+\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2})\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}-I.$
Of this operator, $(\gamma,v)\in\mathbb{R}_{+}\times\mathcal{H}$ is an eigenpair if and only if $w:=\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2}v$ satisfies
$(I+\mathcal{C}_{\textup{pr}}^{1/2}UU^{*}\mathcal{C}_{\textup{pr}}^{1/2})w=(1+\gamma)\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}_{\textup{pos}}^{-1/2}v=(1+\gamma)(\mathcal{C}_{\textup{pr}}^{1/2}\mathcal{C}_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}^{1/2})w.$
Defining $Z:=\mathcal{C}_{\textup{pr}}^{1/2}U$ and using (3b) and (2), the previous equation yields
$(I+ZZ^{*})w=(1+\gamma)(\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}+I)w.$
In this way, we recover the formulation that [30] takes as a starting point to solve the optimization problem. See the proof of [30, Theorem 2.3].
@@ 4.2 Differentiability and minimisers of covariance loss function
$J_{f}(U)=F_{f}(\Lambda(R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})))=F_{f}\circ\Lambda\circ g\ (U).$ (19)
which yields the desired reformulation of the loss as a composition of functions. We search for a solution of 4.4 in the set of stationary points of $J_{f}$. For this, we need to show that $J_{f}$ is Gateaux differentiable. To do so, we use the following result, which states that $g$ and $F_{f}$ Fréchet differentiable and gives an explicit form of the derivatives.

### v3

@@ 4.2 Equivalence to target measures of low-rank Gaussian approximations
    $\mathcal{C}\geq 0$ if and only if $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ and $\mathfrak{X}\geq 0$,
  3. (iii)
    The following are equivalent:
    1. (a)
      $\mathcal{C}\in\mathcal{E}(m_{1},\mathcal{C}_{1})$, with $\mathcal{E}(m_{1},\mathcal{C}_{1})$ defined in (7),
    2. (b)
      $\mathcal{C}>0$ and $\operatorname{ran}{\mathcal{C}^{1/2}}=\operatorname{ran}{\mathcal{C}}_{1}^{1/2}$,
    3. (c)
      $\mathcal{C}\geq 0$ and $\operatorname{ran}{\mathcal{C}^{1/2}}=\operatorname{ran}{\mathcal{C}}_{1}^{1/2}$,
    4. (d)
      $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$, $\mathfrak{X}\geq 0$ and $\operatorname{ran}{\mathcal{C}^{1/2}}=\operatorname{ran}{\mathcal{C}}_{1}^{1/2}$,
    5. (e)
      $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$, $\mathfrak{X}\geq 0$ and $\mathfrak{X}$ is invertible.
  4. (iv)
    Let $\mathcal{C}\geq 0$. Then $\mathfrak{X}=I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}d_{i}^{2}e_{i}\otimes e_{i}$ with $(d_{i}^{2})_{i=1}^{\operatorname{rank}\left({K}\right)}\subset(0,1]$ nonincreasing and $(e_{i})_{i=1}^{\operatorname{rank}\left({K}\right)}$ orthonormal. The equivalent statements of item (iii) hold if and only if $(d_{i}^{2})_{i}\subset(0,1)$. If additionally $\mathcal{C}>0$ and $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}}$, then $(d_{i}^{2})_{i=1}^{\operatorname{rank}\left({K}\right)}\subset(0,1)$ and $(e_{i})_{i=1}^{\operatorname{rank}\left({K}\right)}\subset\operatorname{ran}{\mathcal{C}}_{1}^{1/2}$.
With Lemma 4.5 describing those elements in $\mathcal{E}$ of the form $\mathcal{C_{\textup{pr}}}-KK^{*}$ with $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, we can now characterise the inverses of these elements, and also the inverses of the elements in $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$. As we did for Lemma 4.5, we state the result for low-rank approximations of injective covariances of arbitrary Gaussian measures, rather than only for the prior.
Proposition 4.6.
Let $\mathcal{C},\mathcal{C}_{1}\in L_{1}(\mathcal{H})_{\mathbb{R}}$, $m_{1}\in\mathcal{H}$ and $r\in\mathbb{N}$. Suppose $\mathcal{C}_{1}$ is injective. The following hold:
  1. (i)
    $\mathcal{C}=\mathcal{C}_{1}-KK^{*}$ for $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ and $\mathcal{C}\in\mathcal{E}(m_{1},\mathcal{C}_{1})$ if and only if $\mathcal{C}$ is injective and $\mathcal{C}^{-1}=\mathcal{C}_{1}^{-1/2}(I+ZZ^{*})\mathcal{C}_{1}^{-1/2}$ on $\operatorname{ran}{\mathcal{C}}$ for some $Z\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. In this case, $\operatorname{rank}\left({Z}\right)=\operatorname{rank}\left({K}\right)$.
  2. (ii)
    $\mathcal{C}=\mathcal{C}_{1}-KK^{*}$ for $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, $\mathcal{C}\in\mathcal{E}(m_{1},\mathcal{C}_{1})$ and $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}}$ if and only if $\mathcal{C}$ is injective, $\operatorname{ran}{\mathcal{C}}=\operatorname{ran}{\mathcal{C}_{1}}$ and $\mathcal{C}^{-1}=\mathcal{C}_{1}^{-1}+UU^{*}$ on $\operatorname{ran}{\mathcal{C}_{1}}$ for some $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. In this case, $\operatorname{rank}\left({U}\right)=\operatorname{rank}\left({K}\right)$.
  3. (iii)
    $\mathcal{C}=\mathcal{C}_{1}-KK^{*}$ for $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, $\mathcal{C}>0$ and $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}}$ if and only if $\mathcal{C}$ is injective, $\operatorname{ran}{\mathcal{C}}=\operatorname{ran}{\mathcal{C}_{1}}$ and $\mathcal{C}^{-1}=\mathcal{C}_{1}^{-1}+UU^{*}$ on $\operatorname{ran}{\mathcal{C}_{1}}$ for some $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. In this case, $\operatorname{rank}\left({U}\right)=\operatorname{rank}\left({K}\right)$.
Item (iii) is a slight reformulation of item (ii), and is useful in view of the definition (4) and (5), because with $(m_{1},\mathcal{C}_{1})\leftarrow(0,\mathcal{C}_{\textup{pr}})$ it shows that $\mathscr{C}_{r}^{-1}=\mathscr{P}_{r}$. This fact is summarised in Corollary 4.8(i). Furthermore, a comparison of item (ii) and item (iii) shows that for $\mathcal{C}_{\textup{pr}}-KK^{*}$ with $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, we have the equivalent statements
  1. (i)
    $\mathcal{C}>0$ and $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$, and
  2. (ii)
    $\mathcal{C}\in\mathcal{E}$ and $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$.
Hence, $\mathscr{C}_{r}\subset\mathcal{E}$. This fact is reiterated in Corollary 4.9.
We comment on the difference between the statements in items (i) and (ii). If the equivalent conditions of item (i) hold, then item (i) implies $(I+ZZ^{*})\mathcal{C}_{1}^{-1/2}h\in\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ for any $h\in\operatorname{ran}{\mathcal{C}}$. However, this does not imply that $k_{1}\coloneqq\mathcal{C}_{1}^{-1/2}h$ and $k_{2}\coloneqq ZZ^{*}\mathcal{C}_{1}^{-1/2}h$ lie in $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$, only that their sum $k_{1}+k_{2}$ does. Under the additional condition that $k_{1},k_{2}\in\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$, we may write $\mathcal{C}^{-1}h=\mathcal{C}_{1}^{-1/2}k_{1}+\mathcal{C}_{1}^{-1/2}k_{2}=\mathcal{C}_{1}^{-1}h+\mathcal{C}_{1}^{-1/2}ZZ^{*}\mathcal{C}_{1}^{-1/2}h$. We can formulate the latter as $\mathcal{C}_{1}^{-1}h+UU^{*}h$ for a suitable $U$, as shown in the proof. Thus, to be able to write $\mathcal{C}_{1}^{-1/2}(I+ZZ^{*})\mathcal{C}_{1}^{-1/2}=\mathcal{C}_{1}^{-1}+\mathcal{C}_{1}^{-1/2}ZZ^{*}\mathcal{C}_{1}^{-1/2}$ on all of $\operatorname{ran}{\mathcal{C}}$, as in the formulation of item (ii), one needs to impose restrictions on $\operatorname{ran}{\mathcal{C}}$, and hence on $\operatorname{ran}{K}$, in item (i). As item (ii) shows, the required condition is precisely $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}}_{1}$.
We give an example of $\mathcal{C}=\mathcal{C}_{1}-KK^{*}$ with $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ for which $\mathcal{C}\in\mathcal{E}(m_{1},\mathcal{C}_{1})$ but not $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}}$, which shows that the additional condition $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{1}}$ in item (ii) compared to item (i) is not vacuous. Let $\mathcal{H}$ be infinite-dimensional, so that $\operatorname{ran}{\mathcal{C}_{1}}$ is a proper subspace of $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$. Let $h\in\operatorname{ran}{\mathcal{C}_{1}^{1/2}}\backslash\operatorname{ran}{\mathcal{C}_{1}}$ and define $k\coloneqq\lVert\mathcal{C}_{1}^{-1/2}h\rVert^{-1}h$ so that $z\coloneqq\mathcal{C}_{1}^{-1/2}k$ has unit norm. With $\varphi$ any unit vector in $\mathbb{R}^{r}$, we define the rank-1 operator $K\coloneqq\frac{1}{2}k\otimes\varphi\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. Hence $\mathcal{C}\coloneqq\mathcal{C}_{1}-KK^{*}=\mathcal{C}_{1}-\frac{1}{4}k\otimes k$ satisfies $\operatorname{ran}{K}=\operatorname{span}{\left(k\right)}\subset\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ and $\operatorname{ran}{K}\not\subset\operatorname{ran}{\mathcal{C}_{1}}$. Furthermore, $I-(\mathcal{C}_{1}^{-1/2}K)(\mathcal{C}_{1}^{-1/2}K)^{*}=I-\frac{1}{4}z\otimes z$ is nonnegative and invertible by Lemma A.9 applied with $e_{1}\leftarrow z$, $\delta_{1}\leftarrow-\frac{1}{4}$ and $\delta_{i}\leftarrow 0$ for $i>1$. By the implication (iii)(e)$\Rightarrow$(iii)(a) in Lemma 4.5(iii), $\mathcal{C}\in\mathcal{E}(m_{1},\mathcal{C}_{1})$, which furnishes the desired example.
In our Bayesian context, i.e. setting $(m_{1},\mathcal{C}_{1})\leftarrow(0,\mathcal{C}_{\textup{pr}})$, Lemma 4.5(iii) shows that for all $\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}$ which satisfy $\mathcal{C}\in\mathcal{E}$ and $\mathcal{C}=\mathcal{C}_{\textup{pr}}-KK^{*}$ for some $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, it holds that $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$. However, as Proposition 4.6(i)-(ii) shows, not all $\mathcal{C}\in L_{1}(\mathcal{H})_{\mathbb{R}}$ which satisfy $\mathcal{C}\in\mathcal{E}$ and $\mathcal{C}=\mathcal{C}_{\textup{pr}}-KK^{*}$ for some $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ have associated precision of the form (5), only those for which not only $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ but also $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$ holds. In general, the precisions of covariance operators $\mathcal{C}\in\mathcal{E}$ of the form $\mathcal{C}_{\textup{pr}}-KK^{*}$ are of the form $\mathcal{C}_{\textup{pr}}^{-1/2}(I+ZZ^{*})\mathcal{C}_{\textup{pr}}^{-1/2}$. Of course, if $\dim{\mathcal{H}}<\infty$, then $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}=\mathcal{H}$ and both forms always agree, so that the difference between Proposition 4.6(i)-(ii) disappears.
Since $\operatorname{ran}{\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})^{-1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$, the update (3b) of $\mathcal{C}_{\textup{pr}}$ and item (ii) or item (iii) of Proposition 4.6 with $r\leftarrow n$ and $(m_{1},\mathcal{C}_{1})\leftarrow(0,\mathcal{C}_{\textup{pr}})$ show that $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$. This provides another argument showing that $\operatorname{ran}{\mathcal{C}_{\textup{pr}}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$, besides the explicit computation of [32, Example 6.23].
Remark 4.7 (Choice of approximation class).
A natural generalisation to infinite dimensions of the low-rank approximation classes for covariance and precision considered in the finite-dimensional setting of [31, eqs. (2.4) and (4.1)], is to take $\widetilde{\mathscr{C}}_{r}\coloneqq\{\mathcal{C}_{\textup{pr}}-KK^{*}>0:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$ and $\widetilde{\mathscr{P}}_{r}=\{\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}:\ U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$. Let $\mathscr{C}_{r,\mathcal{E}}\coloneqq\{\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathcal{E}:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$ as in the start of Section 4.2. By the preceding discussion, we have the proper inclusions $\widetilde{\mathscr{P}}_{r}^{-1}\subset\mathscr{C}_{r,\mathcal{E}}$, and by the equivalence (iii)(b)$\Leftrightarrow$(iii)(a) of Lemma 4.5(iii) also the proper inclusion $\mathscr{C}_{r,\mathcal{E}}\subset\widetilde{\mathscr{C}}_{r}$ holds. That is, in general $\widetilde{\mathscr{P}}_{r}^{-1}$ contains strictly fewer covariances than those that maintain equivalence between the resulting approximate and exact posterior, while $\widetilde{\mathscr{C}}_{r}$ constains strictly more. For the loss classes considered in this work which require densities to exist, the set $\widetilde{\mathscr{C}}_{r}$ is thus not suitable. Note that by the definition of $\mathscr{P}_{r}$ in (5), $\widetilde{\mathscr{P}}_{r}=\mathscr{P}_{r}$. The fact that $\widetilde{\mathscr{P}}_{r}^{-1}\subset\mathscr{C}_{r,\mathcal{E}}$ motivates the use of approximation class $\widetilde{\mathscr{P}}_{r}$ in this work. Because the form of the precision updates in $\widetilde{\mathscr{P}}_{r}$ parallels the form in [31, eq. (4.1)] that lies at the core of the approximation procedure of [31], our work can be considered to naturally generalise [31].
Corollary 4.8(i) below shows that $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$ are in one-to-one correspondence by the operation of taking inverses, and can be seen as a generalisation of the finite-dimensional result [31, Lemma A.2] to infinite-dimensional Hilbert spaces. Corollary 4.8(ii) shows that one may solve Problem 4.3 by solving Problem 4.4 and vice versa.
Corollary 4.8.
Let $r\in\mathbb{N}$ and let $\mathscr{C}_{r}$ and $\mathscr{P}_{r}$ be as in (4) and (5) respectively.
  1. (i)
    For every $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ such that $\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathscr{C}_{r}$, there exists $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the same rank as $K$, such that $(\mathcal{C}_{\textup{pr}}-KK^{*})^{-1}=\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\in\mathscr{P}_{r}$. The reverse correspondence also holds: for every $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ such that $\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}\in\mathscr{P}_{r}$, there exists $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the same rank as $U$, such that $(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}=\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathscr{C}_{r}$. In particular, $\mathscr{C}_{r}^{-1}\coloneqq\{\mathcal{C}^{-1}:\ \mathcal{C}\in\mathscr{C}_{r}\}=\mathscr{P}_{r}$ and $\mathscr{P}_{r}^{-1}\coloneqq\{\mathcal{P}^{-1}:\ \mathcal{P}\in\mathscr{P}_{r}\}=\mathscr{C}_{r}$.
  2. (ii)
@@
    An approximation $\mathcal{C}^{\textup{opt}}_{r}\in\mathscr{C}_{r}$ solves Problem 4.3 if and only if $(\mathcal{C}^{\textup{opt}}_{r})^{-1}\in\mathscr{P}_{r}$ solves Problem 4.4. Furthermore, $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{r}^{\textup{opt}})=\mathcal{L}(\mathcal{C}_{\textup{pos}}\|(\mathcal{P}_{r}^{\textup{opt}})^{-1})$.
As discussed after Proposition 4.6 it holds that $\mathscr{C}_{r}\subset\mathcal{E}$. Hence Problem 4.3 is well-defined in the sense that $\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\cdot)$ is finite on $\mathscr{C}_{r}$ for any $\mathcal{L}\in\mathscr{L}$. By Corollary 4.8(ii), it follows that Problem 4.4 is analogously well-defined. These facts are emphasised below.
Corollary 4.9.
It holds that $\mathscr{C}_{r}\subset\mathcal{E}$. Thus, for any $\mathcal{L}\in\mathscr{L}$, the map $\mathcal{C}\mapsto\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{C})$ is finite on $\mathscr{C}_{r}$ and the map $\mathcal{P}\mapsto\mathcal{L}(\mathcal{C}_{\textup{pos}}\|\mathcal{P}^{-1})$ is finite on $\mathscr{P}_{r}$.
4.3 Differentiability and minimisers of covariance loss function
In order to solve Problem 4.4, we formulate it as a minimisation problem over the set of $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. By Corollary 4.8(ii), solving Problem 4.3 is equivalent to solving Problem 4.3. We want to find the minimiser of the function
$J_{f}:\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\rightarrow\mathbb{R},\quad U\mapsto\mathcal{L}_{f}(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1}),$ (17)
for any $f\in\mathscr{F}$ and $\mathcal{L}_{f}\in\mathscr{L}$ defined in (16), which we shall express as a composition of functions. This composition will facilitate the analysis of its differentiability and thereby the identification of its stationary points.
As described in Section 1.5, we denote by $\Lambda$ an eigenvalue map defined on the space of Hilbert–Schmidt operators. Fix an arbitrary $f\in\mathscr{F}$. The restriction of $\Lambda$ to the Hilbert–Schmidt operators with eigenvalues sequence in $(-1,\infty)$ shall be postcomposed with the functions
which are well-defined by Lemma 4.1(i). Note that if $P\in\mathcal{B}(\ell_{2}((-1,\infty)))$ is a permutation, i.e. $((Px)_{i})_{i}=(x_{\pi(i)})_{i}$ for some bijection $\pi$ of $\mathbb{N}$, then for every $x\in\ell^{2}((-1,\infty))$ it holds that $F_{f}(Px)=F_{f}(x)$.
$g(U)=\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}-\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2},$ (19)
where $H$ is the Hessian given in (2). That is, $g(U)$ is a nonnegative, self-adjoint, rank-$r$ update of the negative of the posterior-preconditioned Hessian. The image of $g$ in fact consists of Hilbert–Schmidt operators which can be diagonalised in the Cameron–Martin space, as is shown next. This result also motivates the definition of $g$, as it shows that $g(U)$ has the same eigenvalues as $R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})$.
@@ 4.3 Differentiability and minimisers of covariance loss function
Let $r\in\mathbb{N}$, $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ and $g$ be as in (19). Then $\operatorname{rank}\left({g(U)}\right)\leq r+\operatorname{rank}\left({H}\right)$ and there exists a sequence $(e_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ which forms an ONB of $\mathcal{H}$ and a sequence $(\gamma_{i})_{i}\in\ell^{2}((-1,\infty))$ satisfying $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$. Finally, the eigenvalues of $g(U)$ and $R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})$ agree, counting multiplicities.
@@
$J_{f}(U)=F_{f}(\Lambda(R(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+UU^{*})^{-1})))=F_{f}\circ\Lambda\circ g\ (U),$ (20)
which yields the desired reformulation of the loss as a composition of functions. We use [5, Theorem 12.4.5 (i)] to search for a solution of Problem 4.4 in the set of stationary points of $J_{f}$. For this, we need to show that $J_{f}$ is Gateaux differentiable. To do so, we use the following result, which states that $g$ and $F_{f}$ are Fréchet differentiable and gives an explicit form of the derivatives. Gateaux- and Fréchet derivatives are infinite-dimensional analogs of directional and total derivatives, see for example [17, Section 3.6], [16, Section 1.4] or [5, Section 12.1] for the definition of Gateaux and Fréchet differentiability.